1. The diagonals of a rectangle have lengths 4 + 2x and 6x – 20.
x =

Answer by checkley77(12844)

4 + 2x = 6x – 20
2X-6X=-20-4
-4X=-24
X=-24/-4
X=6
4+2*6=4+12=16 ANS. FOR THE DIAGONAL.
6*6-20=36-20=16 ANS. FOR THE DIAGONAL.